Circular No. 8520

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005bv AND 2005bw Two apparent supernovae have been found on unfiltered CCD images.

SN 2005bv was reported by V. Lipunov, A. Krylov, V. Kornilov, G. Borisov, D. Kuvshinov, A. Belinski, G. Antipov, E. Gorbovskoy, M. Kuznetsov, S. Potanin, N. Tyurina, and V. Vladimirov, Sternberg Astronomical Institute (MASTER 0.36-m robotic telescope; scale 2′′.1/pixel). SN 2005bw was reported by T. Puckett and R. Gagliano (cf. IAUC 8515; 0.60-m automated supernova patrol telescope).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005bv Apr. 20.94 14^{h}24^{m}07.^{s}44 +26^{o}17′50′′.3 16.5 11′′ E, 9′′ S
2005bw Apr. 28.25 13 33 29.73 +33 02 27.8 17.0 8′′.7 E, 4′′.8 S

Additional approximate magnitudes from the respective discoverers: SN 2005bv, 2004 May 7.92, [18.7; 2005 Apr. 28.9, 16.6–16.7. SN 2005bw in UGC 8539, 2002 Apr. 16, [20.0; 2003 Mar. 10, [20.0; 2005 Apr. 29.21, 17.0 (confirming image by T. Crowley, Chiefland, FL, 0.30-m reflector). SN 1999cf also appeared in UGC 8539 (cf. IAUC 7178).

SUPERNOVAE 2005bf AND 2005bq G. Folatelli, N. Morrell, M. Hamuy, and C. Contreras, Las Campanas

Observatory, report that an optical spectrum (range 380–920 nm) of SN 2005bq (cf. IAUC 8515), obtained in the course of the Carnegie Supernova Project on Apr. 19.35 UT with the Las Campanas 2.5-m du Pont telescope (+ WFCCD spectrograph), reveals it to be a type-Ic supernova. The spectrum is similar to that of SN 1983V at 38 days after maximum light, although the Na I 589.3-nm line is particularly strong in the case of SN 2005bq.

M. Hamuy, C. Contreras, S. Gonzalez, and W. Krzeminski, Carnegie Supernova Project, report that the type-Ic SN 2005bf (cf. IAUC 8507, 8509) is undergoing an unusual photometric behavior, brightening from u′ = 18.69 on Apr. 7 to u′ = 17.90 on Apr. 13, and declining to u′ = 18.61 on Apr. 21, after which it has been re-brightening (reaching u′ = 17.92 on Apr. 29). Similar photometric behaviors have been observed in the g′, r′, and i′ bands, although the dip is less pronounced at longer wavelengths. On Apr. 29, g′ = 16.91, r′ = 16.71, and i′ = 16.76. The light curves are posted

at  http://csp1.lco.cl/~cspuser1/images/OPTICAL LIGHT CURVES/

SN05bf.html. Additional photometric and spectroscopic observations are urged.

2005 April 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green